OpenStudy (anonymous):

THE HARDEST QUESTION EVER ! Karl owes $7,464 on a credit card with a 20.4% interest rate compounded monthly. What is the monthly payment he should make in order to pay off this debt in 12 months, assuming he does not charge any more purchases with the card?

OpenStudy (dumbcow):

something everyone should be able to figure out using a financial calculator.....you input present value , num of pay periods, int rate and it will compute your monthly payment here is the math of it if you have to do it by hand set up a present value equation which is a geometric series \[7464 = P(v+v^{2}+...v^{12})\] where P is monthly payment v is present value of 1 dollar payment \[v = \frac{1}{1+i} = \frac{1}{1+(\frac{.204}{12})}\] sum of geometric series is given as: \[S_{12} = \frac{v(1-v^{12})}{1-v}\] solving for P gives: \[P = 7464*\frac{1-v}{v(1-v^{12})}\] \[P \approx 692.85\]

OpenStudy (dumbcow):

i found a nice credit card calculator http://www.financialcalculator.org/credit-cards/payoff-goal-calculator enter the given info and it will tell you monthly payment of 692.85